A kayak can travel 18 miles downstream in 3 hours, while it would take 9 hours to make the same trip upstream. Find the speed of the kayak in still water as well as the speed of the current.
Answer by lwsshak3(11628)

A kayak can travel 18 miles downstream in 3 hours, while it would take 9 hours to make the same trip upstream. Find the speed of the kayak in still water as well as the speed of the current.
***
let x=speed of kayak in still water
let c=speed of current
x+c=speed of kayak downstream
x-c=speed of kayak upstream
travel time*speed=distance
..
3(x+c)=18
9(x-c)=18
..
3x+3c=18
9x-9c=18
..
9x+9c=54(mult. eq by 3)
9x-9c=18
add
18x=72
x=4
3c=18-3x=18-12=6
c=2
speed of kayak in still water=4 mph
speed of current=2 mph